7.) Solve
0.5x + 4 is less than and equal to 1.3 - 3
The solution is {x l x ____ ____}
Find the indicated outputs for f(x) = 5x(2 exponent) - 2x
f(0) =
f(-1) =
f(2) =

Answer by oscargut(2103)

0.5x + 4 <= 1.3 - 3
0.5x+4 <= -1.7
0.5x <= 2.3
x <= 4.6
The solution is {x / x <= 4.6}
Find the indicated outputs for f(x) = 5x^2 - 2x
f(0) = 5(0)^2-2(0)=0
f(-1) = 5(-1)^2-2(-1)=5+2=7
f(2) = 5(2)^2-2(2)=20-4=16